A rod of length L is pivoted from its top end about an axis perpendicular to the rod to form a physical pendulum. The rod has a non-uniform linear mass density λ(y) = A y2 where y is the distance from the pivoted top end and A is a positive constant. (In practice, this could be realized approximately by having the rod actually be conical in shape with a small cone angle.)
(a) What are the (SI) units of A?
(b) Find the period T of the rod’s oscillations (for small oscillations) in terms of L and perhaps, constants (e.g., π, ⅓, g, ...).
1
Expert's answer
2020-02-10T09:14:41-0500

1.) "\u03bb(y) = A y^2"

here dimensional formula of "\lambda" is [M1L-1T0]

and dimensional formula of "y^2" is [M0L2T0]

hence dimensional formula of A = [M1L-3T0]


2.) Here the mass of the rod is variable so we will find out the moment of inertial of the rod about the axis of rotation by taking a small element of mass 'dm' at a 'x' distance from the axis of rotation

"dI = dm\times x^2"

"dI=\lambda dx\times x^2 \newline dI=Ax^2dx\times x^2\newline 
dI=Ax^4dx"

"\int dI=A\intop x^4dx 
\newline
I = A\intop {^L} x^4dx \newline
I=\dfrac{AL^5}{5}"


now calculating the mass of the rod , lets take a small element dm and integrate it

"dm=\lambda dx \newline
\int dm= \int ^L Ax^2dx
m=\dfrac{AL^3}{3}"


Time period is given by "T=2\pi\sqrt{\dfrac{I}{mgd}}" where I is the moment of inertia , m is the mass of the rod and d is distance between the axis of rotation and the end point of the rod

"T=2\pi\sqrt{\dfrac{AL^5/5}{(AL^3/3)\times g\times L}}=2\pi\sqrt{\dfrac{5L}{3g}}"